# Towards Learning Contrast Kinetics with Multi-Condition Latent Diffusion Models

### v1

@@ Abstract
Contrast agents in dynamic contrast enhanced magnetic resonance imaging allow to localize tumors and observe their contrast kinetics, which is essential for cancer characterization and respective treatment decision-making. However, contrast agent administration is not only associated with adverse health risks, but also restricted for patients during pregnancy, and for those with kidney malfunction, or other adverse reactions. With contrast uptake as key biomarker for lesion malignancy, cancer recurrence risk, and treatment response, it becomes pivotal to reduce the dependency on intravenous contrast agent administration. To this end, we propose a multi-conditional latent diffusion model capable of acquisition time-conditioned image synthesis of DCE-MRI temporal sequences. To evaluate medical image synthesis, we additionally propose and validate the Fréchet radiomics distance as an image quality measure based on biomarker variability between synthetic and real imaging data. Our results demonstrate our method’s ability to generate realistic multi-sequence fat-saturated breast DCE-MRI and uncover the emerging potential of deep learning based contrast kinetics simulation. We publicly share our accessible codebase at https://github.com/RichardObi/ccnet.
@@ 3.3 Generation of DCE Sequences from Pre-Contrast Images
CC-Net P1 | Real DCE-P1 | 35.64 | 41.38 | 0.192.070 | 46.9215.40 |
CC-Net P1 | Real DCE-P1 | 39.98 | 42.85 | 0.186.073 | 47.2015.76 |
CC-Net P1 | Real DCE-P1 | 41.55 | 62.41 | 0.200.074 | 49.1714.48 |
CC-Net P1 | Real DCE-P1 | 22.50 | 64.61 | 0.194.072 | 46.1214.21 |
CC-Net P1 | Real DCE-P1 | 45.19 | 60.58 | 0.193.073 | 48.0814.44 |
CC-Net P1 | Real DCE-P1 | 37.70 | 62.21 | 0.192.072 | 45.4813.60 |
@@
In CC-Net, we first initialize a pretrained autoencoder (AE) from stable diffusion (SD), and then (b) proceed to train the denoising U-Net (i.e., LDM) before (c) training the ControlNet, after which we finally (d) run inference on the test set. In each step, we observe several hyperparameters to influence the empirical results. In (a), we note visible differences between the generally accurate breast MRI reconstructions of different pretrained SD AEs. We select the AE from SD 2-1-base over v1-5 and xl-base-1.0. In (b), we notice a high dependence of output quality on scaling factor s with which the AE representation is multiplied before being used in U-Net training. We find s=0.1 to improve results upon the recommended s=0.18215. We further identify a tendency for exploding gradients which was better addressed by clipping the gradient value rather than its norm (e.g., at 15). Both fine-tuning the SD U-Net, but also training it from scratch produced desirable outputs. We use a DDPM [6] noise scheduler with 1000 timesteps, AdamW as optimizer, and batch sizes of 8, 16, and 32, and learning rates of and , for which we obtain similar results. The U-Net is trained for 100 epochs and selected from the epoch with lowest validation loss for further use in (c) and (d). In (c), we follow the hyperparameter setup from (b) with half the batch size and no gradient clipping. In (d), we increase inference speed by using a DDIM [18] scheduler with 200 timesteps without visible performance decrease. We set text guidance scale to 1 as higher scales did not improve output quality, however, increasing the ControlNet guidance weight from 1 to a value in range (1, 1.6] enhanced perceived image quality. In (a)-(c), following AE pretraining, the 224x224 input images are stacked in 3-channels and normalized in range [-1, 1] before applying small-scale intensity and affine augmentations. As observable in Fig. 2(a), the contrast-enhanced tumors account for a large part of the difference between pre- and post-contrast images. Bearing this in mind, we design experiments comparing the image and distribution-wise difference between our synthetic and the real post-contrast images against the difference between corresponding real pre-contrast and real post-contrast images. To quantify this difference, we use our FRD and the common FID [5], LPIPS [24] and mean squared error (MSE) metrics. Obtained results alongside ablations are summarized in Table 1. Overall, CC-Net achieves substantially better results than the pre-contrast baseline across metrics and across post-contrast phases. Surprisingly, providing textual information not necessarily increases the performance of CC-Net. Training the LDM from scratch instead of SD fine-tuning rather decreases performance (e.g., FID). The impact of longer ControlNet training is positive although marginal. Increasing the ControlNet guidance weight can be beneficial, and, in line with qualitative visual analysis, improves FRD and LPIPS. In Fig. 2(b), using CC-Net, we further analyze the pixel intensity distributions in the tumor area across phases aggregated over test cases. Synthetic images noticeably follow the principal trend of real contrast kinetics, despite a scale difference. A similar pattern is observable when taking contrast enhancements outside the tumor area into account by normalizing (dividing) mean tumor intensity by mean intensity of all other (tumor-free) pixels.

### v3

@@ Abstract
Contrast agents in dynamic contrast enhanced magnetic resonance imaging allow to localize tumors and observe their contrast kinetics, which is essential for cancer characterization and respective treatment decision-making. However, contrast agent administration is not only associated with adverse health risks, but also restricted for patients during pregnancy, and for those with kidney malfunction, or other adverse reactions. With contrast uptake as key biomarker for lesion malignancy, cancer recurrence risk, and treatment response, it becomes pivotal to reduce the dependency on intravenous contrast agent administration. To this end, we propose a multi-conditional latent diffusion model capable of acquisition time-conditioned image synthesis of DCE-MRI temporal sequences. To evaluate medical image synthesis, we additionally propose and validate the Fréchet radiomics distance as an image quality measure based on biomarker variability between synthetic and real imaging data. Our results demonstrate our method’s ability to generate realistic multi-sequence fat-saturated breast DCE-MRI and uncover the emerging potential of deep learning based contrast kinetics simulation. We publicly share our accessible codebase at https://github.com/RichardObi/ccnet and provide a user-friendly library for Fréchet radiomics distance calculation at https://pypi.org/project/frd-score.
@@ 3.3 Generation of DCE Sequences from Pre-Contrast Images
CC-NetAny P1 | Real DCE-P1 | 35.64 | 41.38 | 0.192.070 | 46.9215.40 |
CC-NetAny+Txt P1 | Real DCE-P1 | 39.98 | 42.85 | 0.186.073 | 47.2015.76 |
CC-NetAny+Txt+LDM P1 | Real DCE-P1 | 41.55 | 62.41 | 0.200.074 | 49.1714.48 |
CC-NetAny+Txt+LDM+CG1.6 P1 | Real DCE-P1 | 22.50 | 64.61 | 0.194.072 | 46.1214.21 |
CC-NetAny+Txt+LDM+LT P1 | Real DCE-P1 | 45.19 | 60.58 | 0.193.073 | 48.0814.44 |
CC-NetAny+Txt+LDM+CG1.6+LT P1 | Real DCE-P1 | 37.70 | 62.21 | 0.192.072 | 45.4813.60 |
@@
In CC-Net, we first initialize a pretrained autoencoder (AE) from stable diffusion (SD), and then (b) proceed to train the denoising U-Net (i.e., LDM) before (c) training the ControlNet, after which we finally (d) run inference on the test set. In each step, we observe several hyperparameters to influence the empirical results. In (a), we note visible differences between the generally accurate breast MRI reconstructions of different pretrained SD AEs. We select the AE from SD 2-1-base over v1-5 and xl-base-1.0. In (b), we notice a high dependence of output quality on scaling factor s with which the AE representation is multiplied before being used in U-Net training. We find s=0.1 to improve results upon the recommended s=0.18215. We further identify a tendency for exploding gradients which was better addressed by clipping the gradient value rather than its norm (e.g., at 15). Both fine-tuning the SD U-Net, but also training it from scratch produced desirable outputs. We use a DDPM [6] noise scheduler with 1000 timesteps, AdamW as optimizer, and batch sizes of 8, 16, and 32, and learning rates of and , for which we obtain similar results. The U-Net is trained for 100 epochs and selected from the epoch with lowest validation loss for further use in (c) and (d). In (c), we follow the hyperparameter setup from (b) with half the batch size and no gradient clipping. In (d), we increase inference speed by using a DDIM [18] scheduler with 200 timesteps without visible performance decrease. We set text guidance scale to 1 as higher scales did not improve output quality, however, increasing the ControlNet guidance weight from 1 to a value in range (1, 1.6] enhanced perceived image quality. In (a)-(c), following AE pretraining, the 224x224 input images are stacked in 3-channels and normalized in range [-1, 1] before applying small-scale intensity and affine augmentations. As observable in Fig. 3(a), the contrast-enhanced tumors account for a large part of the difference between pre- and post-contrast images. Bearing this in mind, we design experiments comparing the image and distribution-wise difference between our synthetic and the real post-contrast images against the difference between corresponding real pre-contrast and real post-contrast images. To quantify this difference, we use our FRD and the common FID [5], LPIPS [24] and mean squared error (MSE) metrics. Obtained results alongside ablations are summarized in Table 1. Overall, CC-Net achieves substantially better results than the pre-contrast baseline across metrics and across post-contrast phases. Surprisingly, providing textual information not necessarily increases the performance of CC-NetAny. Training the LDM from scratch instead of SD fine-tuning rather decreases performance (e.g., FID). The impact of longer ControlNet training is positive although marginal. Increasing the ControlNet guidance weight can be beneficial, and, in line with qualitative visual analysis, improves FRD and LPIPS. In Fig. 3(b), using CC-NetAny, we further analyze the pixel intensity distributions in the tumor area across phases aggregated over test cases. Synthetic images noticeably follow the principal trend of real contrast kinetics, despite a scale difference. A similar pattern is observable when taking contrast enhancements outside the tumor area into account by normalizing (dividing) mean tumor intensity by mean intensity of all other (tumor-free) pixels.